### v1

# FreeStyle: Free Lunch for Text-guided Style Transfer using Diffusion Models

Abstract
The rapid development of generative diffusion models has significantly advanced the field of style transfer. However, most current style transfer methods based on diffusion models typically involve a slow iterative optimization process, e.g., model fine-tuning and textual inversion of style concept. In this paper, we introduce FreeStyle, an innovative style transfer method built upon a pre-trained large diffusion model, requiring no further optimization. Besides, our method enables style transfer only through a text description of the desired style, eliminating the necessity of style images. Specifically, we propose a dual-stream encoder and single-stream decoder architecture, replacing the conventional U-Net in diffusion models. In the dual-stream encoder, two distinct branches take the content image and style text prompt as inputs, achieving content and style decoupling. In the decoder, we further modulate features from the dual streams based on a given content image and the corresponding style text prompt for precise style transfer. Our experimental results demonstrate high-quality synthesis and fidelity of our method across various content images and style text prompts. The code and more results are available at our project website: https://freestylefreelunch.github.io/.
Figure 1: Illustration depicting process distinctions among Finetune-based, Inversion-based, and our FreeStyle. (a) Fine-tuning the entire model or specific parameters allows embedding a given visual style into the output domain of the generated diffusion model. (b) Embed specific style or content into a new pseudo-word (e.g., ’[C]’) via training set inversion, and use prompts with this pseudo-word to effectively achieve style transfer. (c) In contrast to the above two methods, FreeStyle requires no optimization and utilizes the intrinsic style reconstruction ability of the diffusion model for effective style transfer.
1 Introduction
Image style transfer Jing et al. (2019); Gatys et al. (2015) intends to transfer the natural image into the desired artistic image while preserving the content information. With the recent rapid development of generative diffusion models Rombach et al. (2022); Podell et al. (2023), image style transfer has also witnessed significant advancements. These methods can be broadly classified into two categories: finetuning-based methods Wang et al. (2023); Huang et al. (2022) and inversion-based methods Zhang et al. (2023b); Mokady et al. (2023). The former (depicted in Figure 1 (a)) requires optimizing some or all parameters to degrade the model to generate images of specific styles, while the latter (illustrated in Figure 1 (b)) involves learning the specific style concept as the textual token to guide style-specific generation. Both approaches often require thousands or even more iterations of training, leading to significant computational costs and a slow optimization process.
Large text-guided diffusion models Rombach et al. (2022); Zhang et al. (2023a); Saharia et al. (2022), on the other hand, are typically trained on large-scale datasets of text-image pairs, e.g., LAION dataset Schuhmann et al. (2021), which encompasses various style images and corresponding style text prompts. Consequently, these models Rombach et al. (2022); Podell et al. (2023) inherently possess the generative ability for specific styles. This raises the question: How to leverage the various style generation capabilities of pre-trained text-guided diffusion models for style transfer tasks without additional optimization?
Recent works Alaluf et al. (2023); Hertz et al. (2023) have introduced a cross-image attention mechanism to pre-trained diffusion models, enabling the control of appearance or style transfer without any optimization. However, the use of appearance images or style images as references is still required. In some practical applications, users may not have access to reference images, yet they desire to engage in image transfer based on style text prompts. For instance, users can envision transforming their photos into styles reminiscent of Picasso or Da Vinci without possessing works by these renowned artists. In this paper, we present a novel style transfer approach that requires neither optimization nor style images. Specifically, we propose a novel structure composed of a dual-stream encoder and a single-stream decoder. In this configuration, the dual-stream encoder separately encodes the content image and style text prompt as inputs, extracting features from the corresponding modalities for integration in the decoder. We believe images consist of low-frequency signals controlling image content and high-frequency signals governing image style Seo (2020); Shang et al. (2023). Inspired by FreeU Si et al. (2023), which instantiates two modulation factors to balance low-frequency features from the U-Net’s main backbone and high-frequency features from skip connections to improve the quality of image generation. We modulate the feature maps generated by two distinct encoders using two scaling factors, the first scaling factor regulates the strength of style transfer in the image and the second scaling factor controls the degree of content preservation in the images. Our approach is extremely simple and efficient, requiring only the adjustment of appropriate scaling factors to achieve the transfer of a specific style for any image.
[…]
  • We propose a simple and effective framework based on large text-guided diffusion models, called FreeStyle, which decouples the input of content image and textual input of desired style for specific style transfer without any optimization and style reference.
  • To further balance the preservation of content information and artistic consistency, we propose a novel feature fusion module designed to modulate the features of image content and the corresponding style text prompt.
  • We conducted comprehensive experiments with a wide range of images from various artists and styles. The results show that the art images generated by FreeStyle exhibit accurate style expression and high-quality content-style fusion. Compared to state-of-the-art methods, FreeStyle demonstrates superior and more natural stylization effects.
2 Related Work
2.1 Image Style Transfer
The field of style transfer plays a pivotal role in image processing and computer vision. It has seen a rapid evolution from manual texture synthesis Wang et al. (2004); Zhang et al. (2013) to advanced neural style transfer (NST) Jing et al. (2019); Zhang et al. (2019); Sanakoyeu et al. (2018); Deng et al. (2022); Zhang et al. (2022); He et al. (2023), marking a significant shift from traditional techniques to modern deep learning approaches. Generative Adversarial Networks (GANs) Goodfellow et al. (2014), with impressive image generation capabilities, have been rapidly applied to style transfer tasks Zhu et al. (2017); Karras et al. (2019); Gal et al. (2022), further advancing the development of the field. With the recent rapid development of generative diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020), significant progress has been made in image style transfer. These techniques can be classified into two main categories: finetune-based methods and inversion-based methods. Finetune-based methods Wang et al. (2023); Huang et al. (2022) optimize some or all of the model parameters using extensive style images, embedding their visual style into the model’s output domain. In contrast, inversion-based methods Zhang et al. (2023b); Mokady et al. (2023) embed style or content concepts into special word embeddings using style or content images and achieve style transfer with prompts containing these word embeddings. The aforementioned methods based on diffusion models require style images for training models, leading to a slow optimization process. Recent works Alaluf et al. (2023); Hertz et al. (2023) introduce a cross-image attention mechanism and develop a style transfer method that requires any optimization. However, these methods still rely on style images as references. As a text-guided style transfer method, FreeStyle differs by modulating the latent space features of the diffusion model, leveraging its inherent decoupling ability for style transformation without the need for extra optimization or style reference images.
Figure 2: The overview of our FreeStyle Framework. (a) Model Structure of FreeStyle. Our dual-stream encoder generates the content feature $f_{c}$ guided by the input content image $x_{0}$, and the style feature $f_{s}$ guided by the input style text prompt and noisy image $x_{\sigma}$. In the single-stream decoder, we modulate the content and style features through the feature modulation module. (b) Feature Modulation Module. Our feature modulation module refines style features $f_{s}$ and content features $f_{c}$ separately to ensure accurate style expression and complete content preservation.
2.2 Text-guided Synthesis
GAN-CLS Reed et al. (2016) is the first to achieve text-guided image synthesis of flowers and birds using recurrent neural networks Lipton et al. (2015) and Generative Adversarial Networks Goodfellow et al. (2014). Subsequently, numerous efforts in text-guided image generation Zhang et al. (2017); Xu et al. (2018); Zhu et al. (2019) have propelled rapid development in this field. Benefiting from the introduction of CLIP Radford et al. (2021), GLIDE Nichol et al. (2021) quickly became the first to implement text-guided image generation that conforms to descriptions, following the development of diffusion models. In addition to generating images that match text descriptions, text-guided techniques are now widely used in various tasks such as image editing Wu et al. (2023b); Gal et al. ; Kawar et al. (2023), image restoration Qi et al. (2023); Lin et al. (2023), and video synthesis Esser et al. (2023) etc. Tsu-Jui Fu et al. Fu et al. (2022) argue that traditional style transfer methods, which depend on pre-prepared specific style images, are both inconvenient and creativity-limiting in practical applications. Following this, a new style transfer method that is guided by textual descriptions Kwon and Ye (2022); Patashnik et al. (2021) has been introduced, offering enhanced flexibility and convenience. This not only simplifies complex Artistic creation but also makes advanced image manipulation accessible to a broader audience without the need for specialized graphic design skills. As a result, text-guided image processing is revolutionizing the way we interact with and create visual content.
2.3 Diffusion Models
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) have recently witnessed remarkable advancements, demonstrating their superiority in image generation Nichol et al. (2021); Ramesh et al. (2022); Saharia et al. (2022), video generation Ho et al. (2022); Wu et al. (2023a), 3D generation Watson et al. (2022); Li et al. (2022), etc. Given the impressive generative capabilities of diffusion models, a considerable amount of style transfer works based on diffusion models Zhang et al. (2023b); Wang et al. (2023); Yang et al. (2023); Wu et al. (2023b); Mokady et al. (2023) have been proposed. To further enhance the quality of image synthesis, SDXL Podell et al. (2023) is trained on data with various aspect ratios, utilizing additional attention modules and a larger cross-attention context. Furthermore, FreeU Si et al. (2023) proposed a method to enhance the quality and fidelity of image generation by leveraging the latent capabilities of the U-Net architecture. Similarly, this work utilizes SDXL Podell et al. (2023) as the backbone network to maximize high-quality image synthesis.
3 FreeStyle
3.1 Preliminaries
Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020) involve a forward diffusion process and a reverse denoising process, designed for modeling the mapping between the actual data distribution and the distribution of Gaussian noise. During the forward process, Gaussian noise $\epsilon$ is progressively added to the clean sample $x_{0}$, with the intensity of the added noise $\epsilon$ increasing as $t\in\left[1,2,\dots,T\right]$ increases. The noised image at step $t$ is obtained through the diffusion process:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (1)
where $\epsilon\sim\mathcal{N}\left(0,\mathcal{I}\right)$, $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha_{t}}=\prod\limits_{i=0}^{t}\alpha_{i}$, and $\beta_{t}\in\left(0,1\right)$ is a fixed variance schedule. Conversely, in the denoising process, $x_{T}$ is gradually transformed into the clean image $x_{0}$ by progressively predicting and removing the noise. In DDIM Song et al. (2020), to preserve content information, the deterministic sampling from step $t$ to step $t-1$ can be represented as:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}\left(x_{t}\right)+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{\theta}\left(x_{t},t,c\right).$ (2)
Here, $c$ represents the condition input (e.g., text embedding), and $\epsilon_{\theta}$ denotes the noise prediction network. The predicted denoised image, denoted as $\hat{x}_{0,t}\left(x_{t}\right)$, is obtained by:
$\hat{x}_{0,t}\left(x_{t}\right):=\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}\left(x_{t},t,c\right)}{\sqrt{\bar{\alpha}_{t}}}.$ (3)
To reduce computational demands and achieve high-resolution generation, LDM Rombach et al. (2022) applied diffusion denoising operation on the latent space of a pre-trained encoder.
3.2 Model Structure of FreeStyle
In diffusion models, the U-Net structure is commonly used as the noise prediction network. It consists of an encoder and a decoder, along with skip connections that facilitate information exchange between corresponding layers of the encoder and decoder. Inspired by FreeU Si et al. (2023), which proposes to balance the low-frequency and high-frequency features from the U-Net’s backbone and skip layers, we introduce a novel modulate method of fusing content information and style information applied to style transfer. Figure 2 (a) illustrates the overall structure of FreeStyle, which consists of a dual-stream encoder and a single-stream decoder. The dual-stream encoder in FreeStyle comprises two U-Net encoders with shared parameters, while the single-stream decoder is made up of the U-Net decoder structure. The dual-stream downsampling process can be described separately as follows:
$\begin{cases}f_{s}=&\mathrm{E}\left(x_{\sigma},c\right)\\ f_{c}=&\mathrm{E}\left(x_{0}\right),\\ \end{cases}$ (4)
where $c$ represents the embedding of the style text prompt, and $x_{\sigma}$ denotes the image of the content after $\sigma$ steps of noise addition. $f_{s}$ and $f_{c}$ respectively represent image features that carry style and content information. The denoising process reverses the diffusion process to the predicted clean data $x_{t-1}$ given the nosiy input $x_{t}$:
$p_{\theta}\left(x_{t-1}|x_{t}\right)=$
$\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{0},x_{\sigma},c,t\right),\Sigma_{\theta}\left(x_{0},x_{\sigma},c,t\right)\right).$ (5)
Figure 3: Style transfer results using FreeStyle. Under training-free condition, our method can accurately express its style in images of various categories under various style text prompts, and can achieve a natural fusion of style and content.
Figure 4: The results of style transfer on portraits using FreeStyle. Under the conditions of fine-grained anime style text prompts, the stylized results achieved with FreeStyle still exhibit clear fine-grained style differences and achieve a natural fusion of style and content.
3.3 Feature Modulation Module
FreeU Si et al. (2023) strategically reweights the contributions from U-Net’s skip connections and backbone feature maps, effectively harnessing the strengths of these two components of the U-Net architecture to enhance the quality of the generated images. We believe images consist of low-frequency signals controlling image content and high-frequency signals governing image style. Hence, we implement an effective training-free style transfer by modulating the style feature $f_{s}$ and the content feature $f_{c}$. Different from FreeU, two features that need to be modulated draw from two different inputs, style input $f_{s}$ and content input $f_{c}$.
As shown in Figure 2(b), the content feature $f_{c}$ is generated guided by the noise-free content image $x_{0}$, while the style feature $f_{s}$ is generated guided by the style text prompt $c$ and the noise-added image $x_{\sigma}$. During the upsampling process in U-Net, the features $f_{c}$ primarily influence the semantic expression of the generated result, while the features $f_{s}$ have a greater impact on the high-frequency detail information of the result. Consequently, we engage in special modulation of $f_{s}$ and $f_{c}$ to further activate the intrinsic style reconstruction capability of U-Net. To enhance the semantic characteristics of the feature $f_{c}$, we amplify their variance. Specifically, we apply a weight parameter $b$ greater than 1 to certain dimensions of the feature to expand their variance. We can concisely represent this process as follows:
$f_{c}=concat\left(b\times f_{c}\left[:n\right],f_{c}\left[n:\right]\right).$ (6)
In the formula, the $n$ is used to truncate a portion of the features. On the other hand, to extract style features from the feature $f_{s}$, we believe it is necessary to suppress the low-frequency semantic characteristics while preserving high-frequency details and other style expression information. To achieve this, we first transform the feature $f_{s}$ into frequency domain information using the Fourier transform, and then apply a threshold $r_{\text{thresh}}=1$ to filter out the low-frequency semantic information from the features. Subsequently, we use a weight parameter $s$ greater than 1 to enhance the style information. Finally, we convert the processed frequency domain features back into spatial domain features using the inverse Fourier transform. We can simply denote this process as:
$f_{s}=IFFT\left(\mathcal{F}\left(FFT\left(f_{s}\right)\right)\right),$ (7)
$FFT$ and $IFFT$ represent the Fourier transform and inverse Fourier transform, respectively. The function $\mathcal{F}$ is :
$\mathcal{F}(r)=\left\{\begin{array}[]{ll}s&\text{ if }r<r_{\text{thresh }}\\ 1&\text{ otherwise }\end{array}\right.$ (8)
where $r$ is the radius. Applying the above method, we modulate the $f_{c}$ and the $f_{s}$, and finally concatenate them to feed into the blocks of U-Net decoder.
4 Experiments
In this section, we conduct extensive experiments on images from various domains such as buildings, landscapes, animals, and portrait. By performing qualitative and quantitative comparisons with the state-of-the-art style transfer methods, we validated the robustness and effectiveness of our approach.
4.1 Implementation Details
Since our method is training-free, our method requires no training. Our experiments are conducted on an NVIDIA A100 GPU, with an average sampling time of about 31 seconds for a single image of $1024\times 1024$. All experiments in this section were completed with fixed hyperparameters, set as $n=320$, $s=1$, $b=2.5$, and $\sigma=958$. We use the DDIM sampler to execute a total of 30 sampling steps for each image generation. Our model is based on the SDXL Podell et al. (2023), utilizing its publicly available pre-trained model as the model parameters for inference.
4.2 Experimental Result
Qualitative Results. To verify the robustness and generalization ability of FreeStyle, we conducted numerous style transfer experiments with various styles across different content. Figure 3 presents the style transfer effects of FreeStyle in the domains of buildings, landscapes, animals, etc. The experiments include style transfer in “Chinese Ink”, “Embroidery Art”, “Oil Painting”, “Watercolor Painting”, “Studio Ghibli”, “Cyberpunk”, “Pixel Punk”, “Wasteland” and “Sketching”. We showcase the results of applying style transfer to human portraits using FreeStyle, as in Figure 4. In this figure, we also conducted style transfer experiments with multiple styles, including “Ufotable”, “Studio Ghibli”, “JOJO” and “Illumination Entertainment”. Observations indicate that FreeStyle is capable of providing accurate style information for the style transfer results while almost completely preserving the content information. For instance, the stylized results for “JOJO” maintain the structural information, while reasonably adjusting the image according to the character traits in the “JOJO” anime, like bold outlines, strong lines, and vibrant coloring. This achieves a more natural fusion and expression of both style and content. It is noteworthy that we performed style transfer on images using fine-grained styles from four animation categories in Figure 4. Despite this, FreeStyle is still able to achieve style transfer results with high recognizability and accurate styling.
[…]
Quantitative Comparisons. To better evaluate our method, we employed multiple quantitative metrics for assessment, the results of which are presented in Table 1. For all comparison methods, we utilized their publicly available pretrained parameters for sampling. We performed style transfers on 202 content images including landscapes, architecture, people, and animals, across 10 styles (“Chinese Ink”, “Illumination Entertainment”, “Embroidery Art”, “Graffiti Art”, “Impressionism”, “Oil Painting”, “Watercolor Painting”, “Cyberpunk”, “Studio Ghibli”, “Sketching”), resulting in a total of 2020 stylized images for each method. For the CLIP Score Radford et al. (2021), we calculate the cosine similarity between the CLIP image embeddings and the prompt text embeddings. Using the prompt as a style description, we believe that a higher CLIP Score indicates a more accurate expression of style. The CLIP Aesthetic Score evaluates the quality, aesthetics, and artistic nature of images using a publicly available pre-trained art scoring model. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing.
CLIP Aesthetic Score $\uparrow$ CLIP Score $\uparrow$ Preference $\uparrow$
[…]
FreeStyle (ours) 6.3148 25.615 54.6%
Table 1: Quantitative comparisons with state-of-the-art methods are conducted, using CLIP Aesthetic Score, CLIP Score, and human preferences as our evaluation criteria.
User Study. Additionally, a user study is conducted to evaluate the alignment of stylized images from six methods with human preferences. Forty participants from diverse fields and age groups are invited to evaluate 80 sets of different images, each transferred in various styles using 6 different methods. Participants are asked to select the best stylized image based on considerations of style expression, content preservation, and aesthetics. The “Preference” column in Table 1 reflects the percentage of instances where FreeStyle is ranked first by evaluators, showing its clear superiority over other methods.
[…]
4.3 Ablation Study
Effect of hyperparameters $b$ and $s$. We present the results of ablation experiments conducted on hyperparameters $b$
Figure 6: The ablation study of hyper-parameter $s$ and $b$.
and $s$, as in Figure 6. In FreeStyle, the intensities of content and style information are adjusted by the two hyperparameters $b$ and $s$, respectively. We observed that when the value of $b$ is smaller, the content in the image is more severely disturbed by the style. On the other hand, a larger value of $s$ leads to smoother image textures, while a very small $s$ value creates textures resembling noise. Overall, FreeStyle has relatively low sensitivity to hyperparameters, demonstrating strong robustness. Specifically, we find that its performance is optimal when the hyperparameter $b$ is set to 2.5 and $s$ is set to 1.
Effect of hyperparameter $\sigma$. Figure 7 illustrates the impact of the hyperparameter $\sigma$ on the style transfer effect. The observations indicate that better style transfer are achieved when $\sigma$ exceeds 850, whereas the effect gradually deteriorates as $\sigma$ becomes too small. We believe that a too small $\sigma$ value results in $f_{s}$ containing excessive content information, which significantly disrupts the style information.
Content-Style Disentanglement. To further validate FreeStyle’s ability to disentangle content and style information, we introduced varying degrees of $\rho$ noise into the input $x_{0}$ of the content feature $f_{c}$ to reduce content information input and observed the preservation of content and style information. As shown in Figure 8, with the increase of $\rho$ and hence more noise, the content information in $f_{c}$ gradually decreases while the style feature $f_{s}$ remains unchanged. It is clearly observed that as the value of $\rho$ increases, content information progressively decreases without affecting the expression of style information. When $\rho=999$, content information almost completely disappears, yet the expression of “sketching” style lines and brushstrokes remains observable. This validates FreeStyle’s powerful capability in disentangling content and style information.
Figure 7: The ablation study of hyper-parameter $\sigma$.
Figure 8: An ablation study where varying levels of noise are added to the content image input $x_{0}$ to eliminate content information. (the larger $\rho$, the more noise is introduced)
5 Conclusion
In this work, we introduce FreeStyle, an innovative text-guided style transfer method leveraging pre-trained large text-guided diffusion models. Unlike prior methods, FreeStyle achieves style transfer without additional optimization or the need for reference style images. The framework, featuring a dual-stream encoder and a single-stream decoder, adapts seamlessly to specific style transfer through the adjustment of scaling factors. Despite its extreme simplicity, our method demonstrates superior performance in visual quality, artistic consistency and robust content information preservation across various styles and content images. These results contribute significantly to advancements in the field of training-free style transfer.